OpenStudy (anonymous):

What is true about the solutions of a quadratic equation when the radicand of the quadratic formula is a positive number that is not a perfect square? a. b. c. d.

jimthompson5910 (jim_thompson5910):

If the radicand is not a perfect square, but it's positive, then you have 2 irrational solutions (that are different)

jimthompson5910 (jim_thompson5910):

Ex: x^2 + 6x - 10 has the two solutions \[\Large x = \frac{-6+2\sqrt{19}}{2} \ \text{or} \ x = \frac{-6-2\sqrt{19}}{2}\] since the radicand is 19 (which is a positive non-perfect square)

OpenStudy (anonymous):

a. No real solutions b. Two identical rational solution c. Two different rational solutions d. Two irrational solutions

OpenStudy (anonymous):

So d then !? (:

jimthompson5910 (jim_thompson5910):

you nailed it

jimthompson5910 (jim_thompson5910):

side note: the radicand is known as the discriminant, so if the discriminant is a positive nonperfect square, then you'll also have 2 irrational solutions

OpenStudy (anonymous):

Okie Dokes, thanks !

jimthompson5910 (jim_thompson5910):

yw